# 03F40 (primary), 03D20

### v2

@@ 1. Case 1
    ($r>0$). We use the induction hypothesis on $e$ and $r$ to see that
    $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle m=c^{\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle e}{\uparrow}a+\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle(r-1)\leq d^{\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle e}{\uparrow}a+\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle(r-1)<d^{\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle e}{\uparrow}a+\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle r=\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle n,$
    and since $\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle m=d^{\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle e}{\uparrow}a+\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle(r-1)$, the claim follows.
@@ 2. Case 2
      $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle m$ $=c^{\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle e}{\uparrow}(a-1)+\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle(b^{e}-1)$
      $\leq d^{\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle e}{\uparrow}(a-1)+\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle(b^{e}-1)<d^{\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle e}({\uparrow}a-1)+\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle b^{e}=\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle n,$
@@
      ($a=1$). In this case, the induction hypothesis applied to $b$, $e$, and $b^{e-1}$, and the assumption that $c\leq d$ yields
@@
      $\leq d^{\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle(e-1)}{\uparrow}(b-1)+\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle(b^{e-1}-1)$
      $<d^{\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle e-1}({\uparrow}b-1)+\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle b^{e-1}$
      $\leq d^{\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle e-1}(d-1)+d^{\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle e-1}=d^{\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle e}=\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle n.$
Remark $3.$
Lemma 3.1 does not strictly require $C$ to be a good successor of $B$. In fact, it may even hold when $B$ and $C$ are not base hierarchies! However, the added generality will not be useful to us and it simpler to have this as a ‘blanket’ assumption throughout this section. In the above proof, it allows us to assume that ${\uparrow}^{C}_{B}$ is total. Only Lemma 3.12, fully uses the assumption that $C$ is a good successor, and it is an essential ingredient in our proof of termination.
Next we establish how the upgrade operator interacts with bases. Bellow, ${\rm Base}_{B}$ is as given by Definition 2.3.
@@
  3. 3.
    If $m<n$, $b^{\prime}={\rm Base}_{B}(m)$, and ${\uparrow}m=\langle\begin{smallmatrix}c\\ b^{\prime}\end{smallmatrix}\rangle m$, then $c\leq d$.
Proof 3.5.
1, 2 If $n\in B$, then $n=b$ and $\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle n=\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle b=d\in C$. Moreover, ${\rm Base}_{C}(d)=d$, establishing both claims. If $n\notin B$, then $b<n$. Lemma 3.12 yields
$d=\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle b<\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle n<S_{C}(d).$
This means that ${\uparrow}n\notin C$, since ${\uparrow}n$ is strictly between an element of $C$ and its successor, and moreover ${\rm Base}_{C}({\uparrow}n)=d$ by definition.
3 Towards a contradiction, assume $d<c$. Then, $S_{C}(d)\leq c$, which from the definition of the upgrade operator yields
@@
where the last inequality follows from Lemma 3.12, since $c=\langle\begin{smallmatrix}c\\ b^{\prime}\end{smallmatrix}\rangle{b^{\prime}}\leq\langle\begin{smallmatrix}c\\ b^{\prime}\end{smallmatrix}\rangle(n-1)={\uparrow}(n-1)$. This contradicts Lemma 3.11.
@@
Otherwise, since $n\notin B$, we have that ${\rm Base}_{B}(n)\leq n-1$, from which it follows that ${\rm Base}_{B}(n-1)=b$. We may assume inductively that ${\uparrow}(n-1)=\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle(n-1)$ for the least ${d}\in C$ such that ${\uparrow}b\leq{d}$ and $\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle(n-1)<S_{C}({d})$.
Let $c^{\prime}\geq{\uparrow}b$ be the least element of $C$ such that $\langle\begin{smallmatrix}c^{\prime}\\ b\end{smallmatrix}\rangle n<S_{C}(c^{\prime})$; we must check that $c^{\prime}=c$. By Lemma 3.43 and the induction hypothesis, ${\uparrow}b\leq d\leq c$, so we must indeed have ${\uparrow}b\leq c$. Moreover, $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n<S_{C}(c)$, so $c^{\prime}\leq c$ by minimality of $c^{\prime}$.
In view of Lemma 3.12, $\langle\begin{smallmatrix}c^{\prime}\\ b\end{smallmatrix}\rangle(n-1)<\langle\begin{smallmatrix}c^{\prime}\\ b\end{smallmatrix}\rangle n<S_{C}(c^{\prime})$, yielding $c^{\prime}\geq d$ by minimality of $d$. It follows also from Lemma 3.12 that $\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle(n-1)\leq\langle\begin{smallmatrix}c^{\prime}\\ b\end{smallmatrix}\rangle(n-1)$, and hence ${\uparrow}(n-1)=\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle(n-1)<\langle\begin{smallmatrix}c^{\prime}\\ b\end{smallmatrix}\rangle n$. Since also $\langle\begin{smallmatrix}c^{\prime}\\ b\end{smallmatrix}\rangle n<S_{C}(c^{\prime})$, we conclude that $c^{\prime}\geq c$ by minimality of $c$, and hence equality holds.
@@
Proof 3.9.
The left-to-right direction of the first item is proven by induction on $n$. When $n<b$, the claim follows from $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n={\uparrow}n<{\uparrow}b\leq c$. Otherwise, write $n=_{b}b^{e}a+r$, and note that if $b\mid n$, then $e\geq 1$ and $b\mid r$. The induction hypothesis yields $c\mid\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle r$ and clearly $c\mid c^{\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle e}\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle a=\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle b^{e}a$, so [^1] As per Convention 2.7, $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle b^{e}a=\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle(b^{e}a)$. In the sequel, we apply this convention without mention.
@@
The right-to-left version is similar, but uses Lemma 3.11 to see that every digit of $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n$ is strictly bounded by $c$.
The second item uses Lemma 3.1 to check that the order conditions on $c$-decompositions are satisfied: 3.11 yields $0={\uparrow}0<{\uparrow}a<{\uparrow}b\leq c$ and 3.12 yields $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle r<\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle b^{e}=c^{\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle e}$. The third claim is a corollary of the second by an easy induction.
The fourth item also follows from the second, using the fact that $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n$ yields a $c$-decomposition and comparing the definitions of the left and right sides.
@@
Consider transformations between a base hierarchy $B$ and a good successor $C$. Let $n\in\mathbb{N}$ and $b:={\rm Base}_{B}(n)$. Write $n=ba+r$ with $r<b$ (not necessarily in $b$-decomposition), and let $c\geq{\uparrow}b$. Then,
  1. 1.
    $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n=\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle ab+{\uparrow}r$,
  2. 2.
    ${\uparrow}n={\uparrow}ba+{\uparrow}r$, and
  3. 3.
@@
Proof 3.11.
1 Proceed by induction on $a$. The case for $a=0$ is clear. Otherwise, write $n=_{b}b^{e}x+y$ and note that there is $a^{\prime}$ such that $y=ba^{\prime}+r$ and thus $ba=_{b}b^{e}x+ba^{\prime}$. Then,
$\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n=\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle b^{e}x+\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle(ba^{\prime}+r)\stackrel{{\scriptstyle\text{\sc ih}}}{{=}}\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle b^{e}x+\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle ba^{\prime}+{\uparrow}r=\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle(b^{e}x+ba^{\prime})+{\uparrow}r=\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle ba+{\uparrow}r.$
2 If $a=0$ or $r=0$, there is nothing to prove, so we assume otherwise. Definition 2.5 yields $d,d^{\prime}\in C$ such that ${\uparrow}ba=\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle ba$ and ${\uparrow}n=\langle\begin{smallmatrix}d^{\prime}\\ b\end{smallmatrix}\rangle n$. The claim will follow from showing that $d=d^{\prime}$. Note by Lemma 3.43 that $d\leq d^{\prime}$, so it suffices to check that $d$ satisfies the inequalities required according to Lemma 3.62 for $n$, namely, ${\uparrow}b\leq d$ and $\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle n<S_{C}(d)$.
By Lemmas 3.42 and 3.11, $d\geq{\uparrow}b>{\uparrow}r$. By the definition of a base hierarchy, every element of $C$ divides it successor, so $d\mid S_{C}({\uparrow}ba)$. By Lemma 3.81, $d\mid\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle ba$, hence there are no multiples of $d$ in $(\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle ba,\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle ba+d)$. It follows from 1 that
@@
Thus the desired inequalities hold and conclude that ${\uparrow}n=\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle n={\uparrow}ba+{\uparrow}r$.
3 We have already seen that ${\uparrow}n=\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle n$, so Lemma 3.42 yields ${\rm Base}_{C}({\uparrow}n)=d={\rm Base}_{C}({\uparrow}ba)$.
Next we show how to compute ${\uparrow}n$ when $n$ is a base and $C$ is a good successor. Note that both the upper and lower base are involved in the next lemma; see Definition 2.3.
@@
Let transformations be from $B$ to a good successor $C$ and $b\in B\setminus\{\min B\}$. Let $b^{\prime}$ be the predecessor of $b$. Then, ${\uparrow}b={\uparrow}(b-b^{\prime})+d$, where
@@
Proof 3.13.
Writing $b-1=b-b^{\prime}+r$ with $r=b^{\prime}-1$, we have by Lemma 3.11 that ${\uparrow}r<{\uparrow}b^{\prime}\leq d$. Since $b^{\prime}\mid(b-b^{\prime})$, Lemma 3.102 yields ${\uparrow}(b-1)={\uparrow}(b-b^{\prime})+{\uparrow}r<{\uparrow}(b-b^{\prime})+d$. By Lemma 3.61, ${\uparrow}b=S_{C}({\uparrow}(b-1))$, and since $C$ is a good successor for $B$, this must be the least multiple of $d$ above ${\uparrow}(b-1)$, i.e., ${\uparrow}b^{\prime}={\uparrow}(b-b^{\prime})+d$. Since b1b-1 is not a base, neither is ${\uparrow}(b-1)$ by Lemma 3.41, hence there are no bases in ${\uparrow}(b-1),{\uparrow}b)$, and thus ${\rm base}_{C}({\uparrow}b)=d$.
Finally, we note that the upgrade operator is ‘continuous’: the computation of ${\uparrow}_{B}^{C}n$ only depends on bases below ${\uparrow}_{B}^{C}n$ itself. This can be made precise by restricting $C$ to such values. For $n\in\mathbb{N}$ and a base hierarchy $B$, let $B{\upharpoonright_{n}}=B\cap[0,n]$. Below, note that we are using the upper base as per Definition 2.3.
@@
The proof follows a straightforward induction and is left to the reader. Note that in the third item, there is no restriction on $x$; this is due to the fact that $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle$ only applies the upgrade operator to numbers strictly below $b$.
4 Termination

### v4

@@ 1. Case 1
    ($r>0$). We use the induction hypothesis on $r$ to see that
    $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle m=c^{\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle e}{\uparrow}a+\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle(r-1)<c^{\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle e}{\uparrow}a+\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle r=\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n.$
@@ 2. Case 2
      $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle m$ $=c^{\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle e}{\uparrow}(a-1)+\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle(b^{e}-1)<c^{\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle e}({\uparrow}a-1)+\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle b^{e}=\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n,$
@@
      ($a=1$). In this case, the induction hypothesis applied to $b$, $e$, and $b^{e-1}$, and the assumption that ${\uparrow}b\leq c$ yields
@@
      $<c^{\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle e-1}({\uparrow}b-1)+\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle b^{e-1}$
      $\leq c^{\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle e-1}(c-1)+c^{\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle e-1}=c^{\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle e}=\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n.$
(3) If $n<b$ then $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n={\uparrow}n=\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle n$. Otherwise, write $n=b^{e}a+r$. Induction on $e,r$ yields
$\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n=c^{\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle e}{\uparrow}a+\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle r\leq c^{\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle e}{\uparrow}a+\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle r<d^{\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle e}{\uparrow}a+\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle r=\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle n,$
where the first inequality is by the induction hypothesis and the second uses ${\uparrow}a,\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle e>0$, as can be derived from (1) and (2), given that $a,e>0$.
(4) As above, we have that $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n=c^{\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle e}{\uparrow}a+\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle r$. If $n=b$, then $e=a=1$ and $r=0$, so $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle b=c$ follows. If $n>b$, then $e,a\geq 1$ and either $r>0$ or $\max\{e,a\}>1$, hence $c<\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n$. ∎
Remark 3.2.
Lemma 3.1 does not strictly require $C$ to be a good successor of $B$. In fact, it may even hold in some cases where $B$ and $C$ are not base hierarchies! However, the added generality will not be useful to us and it simpler to have this as a ‘blanket’ assumption throughout this section. For instance, in the above proof, this assumption allows us to assume that ${\uparrow}^{C}_{B}$ is total. Only Lemma 3.7 fully uses the assumption that $C$ is a good successor, but that lemma is indispensable in our proof of termination.
Next, we establish how the upgrade operator interacts with bases. Bellow, ${\rm Base}_{B}$ is as given by Definition 2.3.
@@
  3. (3)
    If $\min B\leq m<n$, $b^{\prime}={\rm Base}_{B}(m)$, and ${\uparrow}m=\langle\begin{smallmatrix}c\\ b^{\prime}\end{smallmatrix}\rangle m$, then $c\leq d$.
Proof.
(1),(2) If $n\in B$, then $n=b$ and $\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle n=\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle b=d\in C$. Moreover, ${\rm Base}_{C}(d)=d$, establishing both claims. If $n\notin B$, then $b<n$. Lemma 3.1(4) yields $d<\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle n<S_{C}(d)$. This means that ${\uparrow}n\notin C$, since ${\uparrow}n$ is strictly between an element of $C$ and its successor, and moreover ${\rm Base}_{C}({\uparrow}n)=d$ by definition.
(3) Towards a contradiction, assume $d<c$. Then, $S_{C}(d)\leq c$, which from the definition of the upgrade operator yields
@@
where the last inequality follows from Lemma 3.1(4), since $c\leq\langle\begin{smallmatrix}c\\ b^{\prime}\end{smallmatrix}\rangle m={\uparrow}m$. This contradicts Lemma 3.1(1). ∎
@@
Otherwise, since $n\notin B$, we have that ${\rm Base}_{B}(n)\leq n-1$, from which it follows that ${\rm Base}_{B}(n-1)=b$. We may assume inductively that ${\uparrow}(n-1)=\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle(n-1)$ for the least ${d}\in C$ such that ${\uparrow}b\leq{d}$ and $\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle(n-1)<S_{C}({d})$ (note that this holds even if n1=bn-1=b).
Let $c^{\prime}\geq{\uparrow}b$ be the least element of $C$ such that $\langle\begin{smallmatrix}c^{\prime}\\ b\end{smallmatrix}\rangle n<S_{C}(c^{\prime})$; we must check that $c^{\prime}=c$. By Lemma 3.3(3) and the induction hypothesis, ${\uparrow}b\leq d\leq c$, so we must indeed have ${\uparrow}b\leq c$. Moreover, $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n<S_{C}(c)$, so $c^{\prime}\leq c$ by minimality of $c^{\prime}$.
In view of Lemma 3.1(2), $\langle\begin{smallmatrix}c^{\prime}\\ b\end{smallmatrix}\rangle(n-1)<\langle\begin{smallmatrix}c^{\prime}\\ b\end{smallmatrix}\rangle n<S_{C}(c^{\prime})$, so $d\leq c^{\prime}$ by minimality of $d$. By Lemma 3.1(3), $\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle(n-1)\leq\langle\begin{smallmatrix}c^{\prime}\\ b\end{smallmatrix}\rangle(n-1)$, and hence ${\uparrow}(n-1)=\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle(n-1)<\langle\begin{smallmatrix}c^{\prime}\\ b\end{smallmatrix}\rangle n$. Since also $\langle\begin{smallmatrix}c^{\prime}\\ b\end{smallmatrix}\rangle n<S_{C}(c^{\prime})$, we conclude that $c^{\prime}\geq c$ by minimality of $c$, and hence equality holds. ∎
@@
Proof.
The left-to-right direction of the first item is proven by induction on $n$. When $n<b$, the claim follows from $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n={\uparrow}n<{\uparrow}b\leq c$. Otherwise, write $n=_{b}b^{e}a+r$, and note that if $b\mid n$, then $e\geq 1$ and $b\mid r$. The induction hypothesis yields $c\mid\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle r$ and clearly $c\mid c^{\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle e}\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle a=\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle b^{e}a$, so [^2] As per Convention 2.7, $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle b^{e}a=\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle(b^{e}a)$.
@@
The right-to-left version is similar, but uses Lemma 3.1(1) to see that every $c$-digit of $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n$ is strictly bounded by $c$.
The second item uses Lemma 3.1 to check that the order conditions on $c$-decompositions are satisfied: 3.1(1) yields $0={\uparrow}0<{\uparrow}a<{\uparrow}b\leq c$ and 3.1(2) yields $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle r<\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle b^{e}=c^{\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle e}$. The third claim is a corollary of the second by an easy induction.
The fourth item also follows from the second, using the fact that $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n$ yields a $c$-decomposition and comparing the definitions of the left and right sides. ∎
@@
Consider transformations between a base hierarchy $B$ and a good successor $C$. Let $n\in\mathbb{N}$ and $b\geq 2$. Write $n=ba+r$ with $r<b$ (not necessarily in $b$-decomposition), and let $c\in C$. Then,
  1. (1)
    $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n=\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle ab+{\uparrow}r$.
  2. (2)
    ${\uparrow}n={\uparrow}ba+{\uparrow}r$.
  3. (3)
    If $\min B\nmid(n+1)$, then ${\uparrow}(n+1)={\uparrow}n+1$.
  4. (4)
@@
Proof.
(1) Proceed by induction on $a$. The case for $a=0$ is clear. Otherwise, write $n=_{b}b^{e}x+y$ and note that there is $a^{\prime}$ such that $y=ba^{\prime}+r$ and thus $ba=_{b}b^{e}x+ba^{\prime}$. Then,
$\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n$ $=\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle b^{e}x+\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle(ba^{\prime}+r)\stackrel{{\scriptstyle\text{\sc ih}}}{{=}}\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle b^{e}x+\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle ba^{\prime}+{\uparrow}r$
$=\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle(b^{e}x+ba^{\prime})+{\uparrow}r=\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle ba+{\uparrow}r.$
(2) If $a=0$ or $r=0$, there is nothing to prove, so we assume otherwise. Definition 2.5 yields $d,d^{\prime}\in C$ such that ${\uparrow}ba=\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle ba$ and ${\uparrow}n=\langle\begin{smallmatrix}d^{\prime}\\ b\end{smallmatrix}\rangle n$. The claim will follow from showing that $d=d^{\prime}$. Note by Lemma 3.3(3) that $d\leq d^{\prime}$, so it suffices to check that $d$ satisfies the inequalities required according to Lemma 3.4(2) for $n$, namely, ${\uparrow}b\leq d$ and $\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle n<S_{C}(d)$.
By Lemmas 3.4 and 3.1(1), $d\geq{\uparrow}b>{\uparrow}r$. By the definition of a base hierarchy, every element of $C$ divides it successor, so $d\mid S_{C}({\uparrow}ba)$. By Lemma 3.5(1), $d\mid\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle ba$, hence there are no multiples of $d$ in $(\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle ba,\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle ba+d)$. It follows from (1) that
@@
Thus the desired inequalities hold and we conclude that ${\uparrow}n=\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle n={\uparrow}ba+{\uparrow}r$.
(3) Let $b=\min B$ and write $n+1=ba+r$ as above, noting that $r>0$. Using (2), ${\uparrow}(n+1)={\uparrow}ba+{\uparrow}r={\uparrow}ba+r$, where the last equality uses $r<\min B$. By the same token, ${\uparrow}n={\uparrow}ba+r-1$, and the claim follows.
(4) We have already seen that ${\uparrow}n=\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle n$ for $d={\rm Base}_{C}({\uparrow}ba)$, so Lemma 3.3(2) yields ${\rm Base}_{C}({\uparrow}n)=d={\rm Base}_{C}({\uparrow}ba)$. ∎
The next lemma shows how to compute ${\uparrow}n$ when $n$ is a base and $C$ is a good successor. Note that both the upper and lower base are involved, as given by Definition 2.3.
@@
Let transformations be from $B$ to a good successor $C$ and let $b\in B\setminus\{\min B\}$. Let $b^{\prime}$ be the predecessor of $b$. Then, ${\uparrow}b={\uparrow}(b-b^{\prime})+d$, where
@@
Proof.
Since $b^{\prime}$ is the predecessor of $b$, $b^{\prime}={\rm Base}_{B}(b-1)$. Writing $b-1=b-b^{\prime}+r$ with $r=b^{\prime}-1$ and letting $d={\rm Base}_{C}({\uparrow}(b-1))$, we have by Lemma 3.1(1) that ${\uparrow}r<{\uparrow}b^{\prime}\leq d$, and by Lemma 3.3(2), ${\uparrow}(b-1)=\langle\begin{smallmatrix}d\\ b^{\prime}\end{smallmatrix}\rangle(b-1)$. Since $b^{\prime}\mid(b-b^{\prime})$, Lemma 3.5(1) yields $d\mid{\uparrow}(b-b^{\prime})$, Lemma 3.6(2) yields ${\uparrow}(b-1)={\uparrow}(b-b^{\prime})+{\uparrow}r<{\uparrow}(b-b^{\prime})+d$, and Lemma 3.6(4) yields $d={\rm Base}_{C}({\uparrow}(b-b^{\prime}))$. By Lemma 3.4(1), ${\uparrow}b=S_{C}({\uparrow}(b-1))$, and since $C$ is a good successor for $B$, this must be the least multiple of $d$ above ${\uparrow}(b-1)$, i.e., ${\uparrow}b^{\prime}={\uparrow}(b-b^{\prime})+d$. Since b1b-1 is not a base, neither is ${\uparrow}(b-1)$ by Lemma 3.3(1), hence there are no bases in ${\uparrow}(b-1),{\uparrow}b)$, and thus ${\rm base}_{C}({\uparrow}b)=d$. ∎
Finally, we note that the upgrade operator is ‘continuous’ in the sense that the computation of ${\uparrow}_{B}^{C}n$ only depends on bases below ${\uparrow}_{B}^{C}n$ itself. This can be made precise by restricting $C$ to such values. For $n\in\mathbb{N}$ and a base hierarchy $B$, let $B{\upharpoonright_{n}}=B\cap[0,n]$.
@@
The proof follows a straightforward induction and is left to the reader. Note that in the second item, there is no restriction on $x$; this is due to the fact that $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle$ only applies the upgrade operator to numbers strictly below $b$.
4. Termination